OpenStudy (anonymous):

9) The width of a rectangle is x inches long. The length of that rectangle is 6 inches less than twice the width. If the perimeter measures 48 inches, what is the measure of the length and width of the rectangle. Show all the work and the algebra you used to solve this problem.

OpenStudy (anonymous):

width = x length = 2*x-5 perimeter = 2*(x + 2x-5) According to question.. 2*(x+2x-5) = 26 => 6x -10 =26 => x = 36/6 = 6 width = 6 inches length = 7 inches.

OpenStudy (anonymous):

since they are two different numbers don't i have to switch them?

OpenStudy (anonymous):

width = x length = 2*x-6 perimeter = 2*(x + 2x-6) According to question.. 2*(x+2x-6) = 48 => 6x -12 =48 => x = 10 width = 10 length =14

OpenStudy (anonymous):

nope everythink is wrong @pkjha3105

OpenStudy (anonymous):

those problmes are easy and peasy

OpenStudy (anonymous):

thts y u hd mistakn to solve it..lol

OpenStudy (anonymous):

no i dont have any mistake man u are mistake the whole problem i have IM

OpenStudy (anonymous):

whre is the mistke?